# Efficient Post-Training Refinement of Latent Reasoning in Large Language Models

### v1

@@ Abstract
Reasoning is a key component of language understanding in Large Language Models. While Chain-of-Thought prompting enhances performance via explicit intermediate steps, it suffers from sufficient token overhead and a fixed reasoning trajectory, preventing step-wise refinement. Recent advances in latent reasoning address these limitations by refining internal reasoning processes directly in the model’s latent space, without producing explicit outputs. However, a key challenge remains: how to effectively update reasoning embeddings during post-training to guide the model toward more accurate solutions. To overcome this challenge, we propose a lightweight post-training framework that refines latent reasoning trajectories using two novel strategies: 1) Contrastive reasoning feedback, which compares reasoning embeddings against strong and weak baselines to infer effective update directions via embedding enhancement; 2) Residual embedding refinement, which stabilizes updates by progressively integrating current and historical gradients, enabling fast yet controlled convergence. Extensive experiments and case studies are conducted on five reasoning benchmarks to demonstrate the effectiveness of the proposed framework. Notably, a +5% accuracy gain on MathQA without additional training. Code and data are publicly available at this link.
Efficient Post-Training Refinement of Latent Reasoning
in Large Language Models
Xinyuan Wang1, Dongjie Wang2, Wangyang Ying1, Haoyue Bai1, Nanxu Gong1, Sixun Dong1, Kunpeng Liu3, Yanjie Fu1 1Arizona State University, 2University of Kansas, 3Portland State University {xwang735, wying4, haoyuebai, ngong6, sixundong, yanjie.fu}@asu.edu wangdongjie@ku.edu, kunpeng@pdx.edu
1 Introduction
@@
Reasoning serves as a fundamental capability in Large Language Models (LLMs), enabling them to comprehend prompts and effectively solve complex tasks. Existing approaches, such as Chain-of-Thought (CoT) Wei et al. (2022b) and ReAct Yao et al. (2023b), guide models toward correct answers by explicitly generating intermediate textual reasoning steps. While these methods have shown effectiveness, they suffer from: 1) the explicit reasoning steps cause substantial token overhead, leading to increased computational cost; 2) the reasoning trajectory becomes fixed once the template is generated, preventing step-by-step refinement during the generation process.
@@
Recent advances have partially addressed them by converting explicit reasoning steps into latent embeddings, enabling latent reasoning in models, such as Coconut Hao et al. (2024). They represent the reasoning state using the LLM’s hidden state (i.e., “continuous thought”) and recursively feed it back into the model in the latent space to enable more effective reasoning. However, there are two critical challenges: 1) the reasoning trajectory in the latent space lacks explicit directional guidance, making it difficult to ensure consistent progression toward more accurate reasoning states; 2) the recursive embedding updates tend to be unstable, especially across multiple reasoning steps, which may compromise both robustness and accuracy. These challenges motivate us to explore how reasoning embeddings can be effectively and efficiently updated during post-training to guide the model toward more accurate solutions.
@@
To this end, we draw inspiration from two complementary lines of research. For the challenge of providing directional guidance in reasoning embedding updates, we are inspired by reinforcement learning from human feedback (RLHF) Ouyang et al. (2022), where learning from relative performance comparisons has demonstrated superior efficiency and effectiveness compared to relying solely on absolute supervision. For the challenge of stabilizing recursive updates, we take inspiration from the success of momentum-based optimization techniques in deep learning Qian (1999), which demonstrate the importance of adaptively integrating historical and current information to achieve smoother and more stable convergence.
@@
Thus, we propose a lightweight post-training framework to refine the latent reasoning embeddings, built upon two novel strategies:
@@
  • Contrastive Reasoning Feedback Search. To infer updated directions in the latent reasoning space, we pass the current reasoning embedding through both a strong and a weak LLM to obtain enhanced embeddings. We derive a contrastive direction by comparing the outputs of strong and weak models, and use its gradient with respect to the current embedding to guide the reasoning embedding update.
@@
  • Residual Embedding Refinement. To ensure stable updates in the latent reasoning space, we blend the current reasoning embedding with its previous state using a residual weighting parameter. This interpolation smooths the transition between steps and prevents abrupt shifts in the reasoning process. As a result, the model achieves more consistent convergence across multi-step latent reasoning.
@@
We empirically evaluate our method through comprehensive experiments and case studies, highlighting its effectiveness, efficiency, and scalability across diverse settings. These experiments demonstrate that our strategies significantly enhance reasoning performance compared to latent-only and explicit token-based reasoning baselines. Notably, on the MathQA task, our approach improves accuracy by over 5% compared to the original latent reasoning method. We further conduct case studies to illustrate how the latent embedding evolves step by step. The results show that the embedding progresses toward more accurate reasoning solutions.
@@
2 Preliminary
2.1 Problem Definition
We focus on complex reasoning tasks where a large language model (LLM) generates a correct answer $y$ from an input question $x$, such as in math word problems, multi-hop question answering, and commonsense reasoning. These tasks typically require multiple inference steps, even if such steps are not explicitly annotated. Formally, the objective is to learn a function $f:x\rightarrow y$, where intermediate cognitive states are latent and only the final answer is observed. While optional intermediate steps can be included during training, they are often unavailable during inference. Building on the latent reasoning framework, we represent each reasoning step as an embedding in a latent space. Our key contribution is to model how to efficiently explore transitions within the reasoning embedding space that lead to accurate final answers. By capturing these latent trajectories, our approach enables the model to reason more effectively, even without explicit supervision over intermediate steps.
2.2 Chain-of-Thought Reasoning and Its Limitations
Chain-of-Thought (CoT) prompting Wei et al. (2022b) and its variants Wang et al. (2022b); Yao et al. (2023b) decompose reasoning into a sequence of intermediate text steps, improving model accuracy and interpretability on complex tasks. However, CoT remains inherently limited:
@@ 2.2 Chain-of-Thought Reasoning and Its Limitations
  • Static trajectory: Once the prompt is fixed, the reasoning path is deterministic, with limited room for correction.
@@
These limitations motivate our shift toward latent-space reasoning. Rather than generating explicit token sequences at each step, we allow the model to evolve its internal reasoning embedding over multiple steps. This latent evolution enables the model to retain richer intermediate information, search for better latent embeddings, and adapt its reasoning process more flexibly, particularly under limited model capacity.
3 Method
@@
Figure 1 shows the overview of our framework. We introduce a general post-training latent refinement framework that enhances latent reasoning models such as Coconut Hao et al. (2024). Our goal is to improve reasoning stability and accuracy by augmenting the latent reasoning process with lightweight, training-free components.
In Coconut, reasoning is performed entirely in latent space without generating token-level intermediate steps, enabling compact and efficient inference. However, the model conducts latent updates in a fixed, feedforward manner, lacking the ability to revise its reasoning path or retain contextual memory across steps—limiting its adaptability when errors occur.
To address these problems, we propose a two-part refinement strategy applied after training:
@@
Both modules operate entirely in latent space through forward passes only, requiring no gradient updates or parameter changes. Applied during inference, they enable accurate and consistent reasoning with minimal computational overhead.
3.1 Latent Reasoning Backbone
We consider a latent reasoning framework in which the model conducts multi-step inference entirely in latent space, without producing token-level intermediate steps. As shown in the top path of Figure 1, the input question $x$ is first encoded into a latent vector $h^{0}$, which is then iteratively updated for $T$ steps using a fixed model block $f$:
@@ 3.1 Latent Reasoning Backbone
where $h^{t}\in\mathbb{R}^{d}$ is the latent embedding at step $t$. After $T$ steps, the final state $h^{T}$ is passed to a decoding head to generate the final answer. This latent-only formulation reduces token overhead and improves inference efficiency, making it particularly well-suited for small or resource-constrained models. This backbone structure forms the basis of our reasoning framework. In our implementation, we adopt Coconut Hao et al. (2024) as the underlying latent reasoning model, where the decoder block $f$ is derived from a pre-trained language model and kept fixed during inference.
@@
    Trajectory stability: Without memory-preserving connections, the latent trajectory may drift or collapse over time.
@@
3.2 Post-Training Latent Reasoning Refinement
@@
3.2.1 Contrastive Reasoning Feedback Search
@@
As shown in Figure 2, we compare outputs from two models of different quality at each reasoning step $t$: A weaker model (“bad” model like earlier checkpoint), producing latent output $h^{t}_{\text{bad}}$; A stronger model (“good” model like later checkpoint), producing output $h^{t}_{\text{good}}$.
@@
$h^{t}_{\text{updated}}=h^{t}+\eta\cdot\nabla_{h^{t}}\big{[}$ $\text{MSE}(h^{t},h^{t}_{\text{good}})$ (2)
$-\text{MSE}(h^{t},h^{t}_{\text{bad}})\big{]}.$
@@
This contrastive search provides a self-correction mechanism during reasoning. It helps the model adjust its latent trajectory without relying on external feedback or additional samples. This mimics the role of conflict monitoring and adjustment of the ACC in human reasoning Botvinick et al. (2001).
Coconut does not have any correction mechanism. If the reasoning goes off track, it cannot adjust or recover. CoT Wei et al. (2022b) and Tree-of-Thoughts Yao et al. (2023a) use external sampling or search to fix errors, but they rely on generating intermediate text. In contrast, our method updates the latent state directly using internal feedback from stronger models. This allows for efficient and flexible reasoning without relying on token-level outputs or any training.
3.2.2 Residual Embedding Refinement
At each step, we update the latent state by preserving useful information from the previous step. Instead of directly replacing the latent with the new output $f(h^{t-1})$, we blend it with the previous state $h^{t-1}$ using a fixed-weight residual connection:
@@ 3.2.2 Residual Embedding Refinement
where $\alpha$ is the memory rate. It controls how much of the previous state is kept. We use a fixed value and do not train this parameter due to the train-free setting. This design is inspired by residual networks He et al. (2016) and resembles the working memory mechanism of the human brain Koechlin et al. (2003). It allows the model to accumulate reasoning context over steps and prevents semantic drift. Without this refinement, the latent state may lose important early signals, which leads to an unstable reasoning trajectory.
@@
4 Experimental Results
We conduct experiments to evaluate whether our latent reasoning framework improves reasoning accuracy, supports training-free deployment, and operates efficiently across diverse tasks and models. Specifically, our experiments are structured to answer the following key questions: Q1: Can our method improve reasoning performance with minimal cost, using only training-free post-processing? Q2: Compared to full model retraining, can our post-training refinement achieve similar improvements with lower resource usage? Q3: How important are the two components—residual refinement and contrastive latent search—in contributing to performance improvement? Q4: Can the latent update mechanism consistently steer the model toward more accurate predictions in specific reasoning instances? Q5: How robust is our method to variations in key hyperparameters such as memory update rate and latent search step size? Q6: Can our framework generalize well across diverse language model architectures and parameter sizes?
4.1 Experimental Setup
Datasets. We evaluate our method on five representative benchmarks covering math, commonsense, and multi-hop reasoning: GSM8K Cobbe et al. (2021), MathQA Amini et al. (2019), AQUA-RAT Ling et al. (2017), StrategyQA Geva et al. (2021), and ProsQA Hao et al. (2024). Dataset descriptions are provided in Appendix B.
Models. We exploit three well-recolonized open-source language models: GPT-2 (117M), Qwen-2.5 1.5B, and LLaMA-3.2 3B. For contrastive search, we utilize checkpoints from different training stages as “good” or “bad” references.
Baselines. We compare our method with the following approaches: (1) No-CoT: directly trains GPT-2 Radford et al. (2019) to generate the final answer without any intermediate reasoning steps; (2)Chain-of-Thought (CoT) Wei et al. (2022b): standard step-by-step natural language reasoning approach; (3) Coconut Hao et al. (2024): latent reasoning without search or refinement; (4) Ours: latent reasoning with contrastive search and residual refinement.
Evaluation. We report exact-match accuracy, averaged over 3 random seeds. No fine-tuning is used; our method improves reasoning purely through forward latent-space updates. Full implementation details are provided in Appendix D.
4.2 Q1: Overall Reasoning Performance
@@
The results highlight that: (1) Although CoT produces explicit thought steps, it suffers from verbosity and error accumulation, especially in non-math tasks. (2) Coconut improves reasoning by operating in a compact latent space, but lacks error correction and stable refinement. (3) Our method combines both advantages through residual preservation and feedback-driven search, thus, result into more reliable and generalizable reasoning, especially in settings where intermediate steps are implicit or hard to verbalize.
@@
4.3 Q2: Training vs. Inference Performance
To answer Q2, we compare our train-free method with other training strategies to examine reasoning accuracy and resource usages under the ProsQA dataset. We train a base model using Coconut for 30 epochs, then evaluate four settings: (1) using the original Coconut model, (2) applying our latent reasoning only during inference, (3) continuing training for 10 more epochs with our latent reasoning enabled, and (4) applying our latent reasoning during both training and inference.
Figure 4 shows that our method achieves the best accuracy (+4.47%) when applied only during inference, with minimal overhead (24 seconds, 31.23GB memory). In contrast, continuing training with latent reasoning adds significant time (54+ minutes) and memory cost (39.04GB), yet leads to smaller accuracy gains (+1.63%). Using the method in both training and inference further increases the cost but gives almost no improvement. These results highlight the efficiency of our strategy: without any backpropagation or weight updates, our inference-only setup improves performance while saving training time and GPU resources and without the need of further training.
@@
4.4 Q3: Study of Contrastive Search and Residual Refinement
To answer Q3, we remove contrastive search and residual refinement one by one to understand the role of each technical component.
@@ 4.4 Q3: Study of Contrastive Search and Residual Refinement
Table 1 highlights the importance of both residual refinement and contrastive search in our framework. Compared with the baseline Coconut with only latent thoughts, our method of incorporating residual connections yields an +4.63% improvement in accuracy. This observation demonstrates that preserving and gradually refining previous latent states help to stabilize reasoning trajectories. It aligns with the working memory mechanism in human prefrontal cortex, where ongoing cognitive representations are maintained and adjusted over time. Only incorporating contrastive search leads to a gain of +4.03%. This observation shows that self-correction based on improving directions, which mimic the role of the anterior cingulate cortex (ACC) in conflict detection, enables the model to recover from suboptimal reasoning directions. Finally, integrating both residential connections and contrastive search can result in the best performance (+5.10%). This observation shows that stable memory evolution and dynamic search together form a lightweight and training-free latent reasoning mechanism that maintains contexts, detects errors, and refines internal representations without relying on explicit intermediate language.
4.5 Q4: A Step-wise Case Study
@@
Insights. The case study illustrates how our method enables internal latent correction before decoding. Rather than relying on external tokens or explicit logic, the model self-adjusts in latent space, which reflects how humans reconsider their thoughts before answering. We provide three more examples in Appendix G.
4.6 Q5: Study of Hyperparameter Sensitivity
@@
Figure 6(a) shows the heatmap of accuracy on MathQA over memory rates and search steps, where red indicates high accuracy and green indicates low accuracy. Overall, the reasoning accuracy increases when the memory rate increases, but is less affected by the search step length. This suggests that mathematical problem solving benefits more from stable accumulation of prior steps and residual refinement (i.e., memory rate). A stronger memory allows the model to preserve intermediate reasoning steps for better reasoning. Figure 6(b) shows the opposite trend. AQUA is less sensitive overall but is more affected by the search step size. This suggests that commonsense QA relies more on adaptive correction than long-term memory, so flexible updates in latent space have a larger impact than memory accumulation.
4.7 Q6: Generalization Across Backbones
To answer Q6, we evaluate the generalization of our framework on different LLM backbones across architectures and sizes: GPT-2 (117M), Qwen-2.5 1.5B, and LLaMA-3.2 1B models. Table 2 shows that our method on top of the three backbone models reach over 40% accuracy on MathQA. This suggests that our latent refinement strategy can be applied to both small and medium-sized models without any changes to model architecture. Our method adds little overhead in terms of resource usage. Meanwhile, training time and memory usage increase with model size as expected, but inference remains efficient. All models finish complete in under 7 minutes with less than 24 GB of memory. Our method is simple and general. It can be used on top of any existing language model to improve reasoning performance, even when training resources are limited. It does not require retraining or modification of model weights, which makes it easy to adopt in different settings.
@@ 4.7 Q6: Generalization Across Backbones
4.8 Study of Token Efficiency
Latent reasoning avoids generating intermediate natural language steps (i.e., thoughts), thus, reduce output token number. We compare the average number of generated tokens between Chain-of-Thought (CoT) prompting and our latent reasoning method on the MathQA and AQUA datasets. Table 3 shows that latent method reduces token usage by over 92% on both datasets. This is because latent steps are performed in a latent space and don’t produce textual outputs at each reasoning step. These results suggest that latent reasoning is not only more compact but also more cost-efficient during inference. This makes it particularly suitable for deployment in resource-constrained settings or large-scale usage.
@@ 4.8 Study of Token Efficiency
5 Related Work
5.1 Chain of Thought Reasoning

### v2

@@ Abstract
Reasoning is a key component of language understanding in Large Language Models. While Chain-of-Thought prompting enhances performance via explicit intermediate steps, it suffers from sufficient token overhead and a fixed reasoning trajectory, preventing step-wise refinement. Recent advances in latent reasoning address these limitations by refining internal reasoning processes directly in the model’s latent space, without producing explicit outputs. However, a key challenge remains: how to effectively update reasoning embeddings during post-training to guide the model toward more accurate solutions. To overcome this challenge, we propose a lightweight post-training framework that refines latent reasoning trajectories using two novel strategies: (1) Contrastive reasoning feedback, which compares reasoning embeddings against strong and weak baselines to infer effective update directions via embedding enhancement; (2) Residual embedding refinement, which stabilizes updates by progressively integrating current and historical gradients, enabling fast yet controlled convergence. Extensive experiments and case studies are conducted on five reasoning benchmarks to demonstrate the effectiveness of the proposed framework. Notably, a +5% accuracy gain on MathQA without additional training.
Codehttps://github.com/anord-wang/Lateng-Reasoning
Introduction
@@
Reasoning serves as a fundamental capability in Large Language Models (LLMs), enabling them to comprehend prompts and effectively solve complex tasks. Existing approaches, such as Chain-of-Thought (CoT) (wei2022chain) and ReAct (yao2023react), guide models toward correct answers by explicitly generating intermediate textual reasoning steps. While these methods have shown effectiveness, they suffer from: (1) the explicit reasoning steps cause substantial token overhead, leading to increased computational cost; (2) the reasoning trajectory becomes fixed once the template is generated, preventing step-by-step refinement during the generation process.
@@
Recent advances have partially addressed them by converting explicit reasoning steps into latent embeddings, enabling latent reasoning in models, such as Coconut (hao2024training). They represent the reasoning state using the LLM’s hidden state (i.e., “continuous thought”) and recursively feed it back into the model in the latent space to enable more effective reasoning. However, there are two critical challenges: (1) the reasoning trajectory in the latent space lacks explicit directional guidance, making it difficult to ensure consistent progression toward more accurate reasoning states; (2) the recursive embedding updates tend to be unstable, especially across multiple reasoning steps, which may compromise both robustness and accuracy. These challenges motivate us to explore how reasoning embeddings can be effectively and efficiently updated during post-training to guide the model toward more accurate solutions.
@@
To this end, we draw inspiration from two complementary lines of research. For the challenge of providing directional guidance in reasoning embedding updates, we are inspired by reinforcement learning from human feedback (RLHF) (ouyang2022training), where learning from relative performance comparisons has demonstrated superior efficiency and effectiveness compared to relying solely on absolute supervision. For the challenge of stabilizing recursive updates, we take inspiration from the success of momentum-based optimization techniques in deep learning (qian1999momentum), which demonstrate the importance of adaptively integrating historical and current information to achieve smoother and more stable convergence.
@@
Thus, we propose a lightweight post-training framework to refine the intermediate latent reasoning embeddings, built upon two novel strategies:
@@
  • Contrastive Reasoning Feedback Search. To infer updated directions in the latent reasoning space, we pass the current reasoning embedding through both a strong and a weak model to obtain enhanced embeddings. We derive a contrastive direction by comparing the outputs of strong and weak models, and use its gradient concerning the current embedding to guide the embedding update. The strong model is only better relative to the weak one, and does not limit the performance of the final method.
@@
  • Residual Embedding Refinement. To ensure stable updates in the latent space, we blend the current reasoning embedding with its previous state using a residual weighting strategy. This interpolation smooths the transition between steps and prevents abrupt shifts in the reasoning process. As a result, the model achieves more consistent convergence across multi-step latent reasoning.
@@
We evaluate our method through comprehensive experiments and case studies, highlighting its effectiveness, efficiency, and scalability across diverse settings. These experiments demonstrate that our strategies significantly enhance reasoning performance compared to latent-only and explicit token-based reasoning baselines. Notably, on the MathQA task, our approach improves accuracy by over 5% compared to the original latent reasoning method. We further conduct case studies to illustrate how the latent embedding evolves step by step. The results show that the embedding progresses toward more accurate reasoning solutions.
@@
Preliminary
Problem Definition
We focus on complex reasoning tasks where a large language model (LLM) aims to generate a correct answer $y$ from an input question $x$, such as in math word problems, multi-hop question answering, and commonsense reasoning. These tasks typically require multiple inference steps, even if such steps are not explicitly annotated. Formally, the objective is to learn a function $f:x\rightarrow y$, where intermediate cognitive states are latent and only the final answer is observed. While optional intermediate steps can be included during training, they are often unavailable during inference. Building on the latent reasoning framework, we represent each reasoning step as an embedding in a latent space. Our key contribution is to model how to efficiently explore transitions within the reasoning embedding space that lead to accurate final answers. By capturing these latent trajectories, our approach enables the model to reason more effectively, even without explicit supervision over intermediate steps.
Chain-of-Thought Reasoning and Its Limitations
Chain-of-Thought (CoT) prompting (wei2022chain) and its variants (wang2022self; yao2023react) decompose reasoning into a sequence of intermediate text steps, improving model accuracy and interpretability on complex tasks. However, CoT remains inherently limited:
@@
  • Static trajectory: Once the Chain-of-Thought prompt is fixed, the reasoning path is deterministic, with limited room for correction.
@@
These limitations motivate our shift toward latent space reasoning. Rather than generating explicit token sequences at each step, we allow the model to evolve its internal embedding over multiple steps. This latent evolution enables the model to retain richer intermediate information, search for better latent embeddings, and adapt its reasoning process more flexibly, particularly under limited model capacity.
Method
Figure 1 shows the overview of our framework. We introduce a general post-training latent refinement framework that enhances latent reasoning models such as Coconut (hao2024training). Our goal is to improve reasoning stability and accuracy by augmenting the latent reasoning process with lightweight, training-free components.
In Coconut, reasoning is performed entirely in latent space without generating token-level intermediate steps, enabling compact and efficient inference. However, the model conducts latent updates in a fixed, feedforward manner. This end-to-end process and training cannot revise its reasoning path or retain contextual memory across steps during inference time, which limits its adaptability when errors occur. Adding such a correction capability during training would require large amounts of data and extensive optimization, making it expensive and less practical.
@@
To address these problems, we propose a two-part refinement strategy applied in the post-training stage:
@@
This strategy is lightweight and training-free. Both modules operate entirely in latent space using forward passes only, without any gradient updates or parameter changes. It allows Coconut to adjust its reasoning during inference by following a contrastive direction, while the residual update prevents overly large changes, helping to retain useful information. Applied at inference time, these components enable more accurate and stable reasoning with minimal computational overhead.
Latent Reasoning Backbone
We consider a latent reasoning framework in which the model conducts multi-step inference entirely in latent space, without producing token-level intermediate steps. As shown in the top path of Figure 1, the input question $x\in\mathbb{R}^{L}$ is first encoded into a latent embedding $h^{0}$, which is then iteratively updated for $T$ steps using a fixed model block $f$:
@@
where $h^{t}\in\mathbb{R}^{L\times d}$ is the latent embedding at step $t$. After $T$ steps, the final state $h^{T}$ is passed to a decoding head to generate the final answer. This latent-only formulation reduces token overhead and improves inference efficiency, making it particularly well-suited for small or resource-constrained models. This backbone structure forms the basis of our reasoning framework. In our implementation, we adopt Coconut (hao2024training) as the underlying latent reasoning model, where the decoder block $f$ is derived from a pre-trained language model and kept fixed during inference.
@@
    Trajectory stability: In multi-step reasoning, the latent trajectory may drift without memory-preserving connections, leading to unstable or inconsistent reasoning.
@@
Post-Training Latent Reasoning Refinement
@@
Contrastive Reasoning Feedback Search
@@
As shown in Figure 2, we compare the outputs from two models of different quality at each reasoning step $t$: a weaker model (“bad” model, such as an early checkpoint) producing latent output $h^{t}_{\text{bad}}$, and a stronger model (“good” model, such as a later checkpoint) producing output $h^{t}_{\text{good}}$.
These two models are not the final model used for latent reasoning inference (e.g., Coconut), but are intermediate checkpoints saved during the training of a standard CoT model. The distinction between “good” and “bad” is relative and only exists for the purpose of creating a directional signal in latent space. The goal is to identify a better direction to refine the current reasoning state, not to imitate or match the stronger model’s behavior.
Importantly, the “good” model used in this comparison is not stronger than our final model. In fact, both “good” and “bad” models are weaker than Coconut. Their role is solely to help define an updated direction based on contrast, rather than to serve as a performance reference or target.
@@
$h^{t}_{\text{updated}}=h^{t}+\eta\cdot\nabla_{h^{t}}\big[$ $\text{MSE}(h^{t}_{\text{good}},h^{t})$ (2)
$-\text{MSE}(h^{t}_{\text{bad}},h^{t})\big].$
@@
This contrastive search provides a self-correction mechanism during reasoning. It helps the model adjust its latent trajectory without relying on external feedback or additional samples. This mimics the role of conflict monitoring and ACC adjustment in human reasoning (botvinick2001conflict).
Currently, Coconut does not have any correction mechanism. If the reasoning goes off track, it cannot adjust or recover. CoT (wei2022chain) and Tree-of-Thoughts (yao2023tree) use external sampling or search to fix errors, but they rely on generating intermediate text. In contrast, our method updates the latent state directly using internal feedback from contrastive direction. This is efficient and flexible, without relying on token-level outputs or any training.
Residual Embedding Refinement
To make the latent embeddings evolve stably, we add a residual mechanism when updating the embeddings. At each step, we update the latent state by preserving useful information from the previous step. Instead of directly replacing the latent with the new output $f(h^{t-1})$, we blend it with the previous state $h^{t-1}$ using a fixed-weight residual connection:
@@
where $\alpha$ is the memory rate. It controls how much of the previous state is kept. We use a fixed value and do not train this parameter due to the train-free setting. This design is inspired by residual networks (he2016deep) and resembles the working memory mechanism of the human brain (koechlin2003architecture). It allows the model to accumulate reasoning context over steps and prevents semantic drift. Without this refinement, the latent state may lose important early signals, which leads to an unstable reasoning trajectory.
In training-free settings, we cannot rely on backward gradients to correct mistakes, so it becomes more important to preserve helpful signals across steps. The residual update plays this role by softly carrying forward past reasoning information, making the trajectory more stable. While the contrastive reasoning search explores better directions in latent space, the residual refinement helps maintain stability during multi-step updates.
@@
Experimental Results
We conduct experiments to evaluate whether our latent reasoning framework improves reasoning accuracy, supports training-free deployment, and operates efficiently across diverse tasks and models. Specifically, our experiments are structured to answer the following key questions:
Q1: Can our method improve reasoning performance with minimal cost, using only training-free post-processing?
Q2: Compared to fully retraining, can our post-training refinement achieve improvements with lower resource usage?
Q3: How important are the two components—residual refinement and contrastive latent search—in contributing to performance improvement?
Q4: Can the latent update mechanism consistently steer the model toward more accurate predictions in specific reasoning instances?
Q5: How robust is our method to changes in hyperparameters like memory update rate and latent search step length?
Q6: Can our framework generalize well across diverse language model architectures and parameter sizes?
Experimental Setup
Datasets. We evaluate our method on five representative benchmarks covering math, commonsense, and multi-hop reasoning: GSM8K (cobbe2021training), MathQA (amini2019mathqa), AQUA-RAT (ling2017program), StrategyQA (geva2021did), and ProsQA (hao2024training). Dataset descriptions are provided in Appendix C.
Models. We use GPT-2 base (117M) to conduct the majority of experiments. We also try two other well-recolonized open-source language models: Qwen-2.5 1.5B and LLaMA-3.2 3B. For contrastive search, we utilize checkpoints from different training stages in the CoT method as “good” or “bad” references.
Baselines. We compare our method with the following approaches: (1) No-CoT: directly trains GPT-2 (radford2019language) to generate the final answer without any intermediate reasoning steps; (2)Chain-of-Thought (CoT) (wei2022chain): standard step-by-step natural language reasoning approach; (3) Coconut (hao2024training): latent reasoning without search or refinement; (4) Ours: latent reasoning with contrastive search and residual refinement.
Evaluation. We report exact match precision, averaged on 3 random seeds. No fine-tuning is used; our method improves reasoning only through forward latent-space updates. Full implementation details are provided in Appendix E.
Q1: Overall Reasoning Performance
@@
The results highlight that: (1) Although CoT produces explicit thought steps, it suffers from verbosity and error accumulation, especially in non-math tasks. (2) Coconut improves reasoning by operating in a compact latent space, but lacks error correction and stable refinement. (3) Our method combines both advantages through residual preservation and feedback-driven search, thus resulting into more reliable and generalizable reasoning, especially in settings where intermediate steps are implicit or hard to verbalize.
@@
Q2: Training vs. Inference Performance
To answer Q2, we compare our train-free method with other training strategies to examine the accuracy of reasoning and the usage of resources in the ProsQA dataset. We train a base Coconut model for 30 epochs, then evaluate four settings: (1) using the original Coconut model, (2) only applying the latent reasoning during inference, (3) continuing training for 10 more epochs with our latent reasoning, and (4) applying our latent reasoning during both training and inference.
Figure 4 shows that our method achieves the best accuracy (+4.47%) when applied only during inference, with minimal overhead (24 seconds, 31.23GB memory). In contrast, continuing training with latent reasoning adds significant time (54+ minutes) and memory cost (39.04GB), yet leads to smaller accuracy gains (+1.63%).
Using the method in both training and inference further increases the cost but gives almost no improvement. This result may seem counter-intuitive at first. But we observed the same pattern on three datasets, so we believe it is consistent. Our understanding is: during training, the refinement only affects the forward pass. It does not go into the backward gradient, so the model may not fully learn how to use it. With limited training data, this extra signal may even confuse the model or hurt convergence. Then, during inference, if we apply refinement again, the model receives a second adjustment. This may “overshoot” the correct direction and make the reasoning worse. That could explain why training + inference gives smaller gains than inference-only.
These results highlight the efficiency of our strategy: without any backpropagation or weight updates, our inference-only setup improves performance while saving training time and GPU resources, and without the need for further training.
@@
Q3: Study of Contrastive Search and Residual Refinement
To answer Q3, we remove the the contrastive search and residual refinement one by one to understand the role of each technical component.
@@
Table 1 highlights the importance of residual refinement and contrastive search in our framework. Compared to the Coconut baseline with only latent thoughts, incorporating residual connections yields a +4.63% improvement in accuracy. This suggests that preserving and gradually refining previous latent states helps stabilize reasoning trajectories across steps. Using only contrastive search leads to a +4.03% gain, showing that updating the current latent state with better directions enables the model to recover from suboptimal reasoning. Finally, integrating both residential connections and contrastive search results in the best performance (+5.10%). This shows that stable memory evolution and dynamic search together form a lightweight and training-free latent reasoning mechanism that maintains contexts, detects errors, and refines internal representations without relying on explicit intermediate language tokens.
Q4: A Step-wise Case Study
@@
Insights. The case study illustrates how our method enables internal latent correction before decoding. Rather than relying on external tokens or explicit logic, the model self-adjusts in latent space, which reflects how humans reconsider their thoughts before answering. We provide three more examples in Appendix H.
Q5: Study of Hyperparameter Sensitivity
@@
Figure 6(a) shows the heat map of accuracy in MathQA on memory rates and search steps, where red indicates high accuracy and green indicates low accuracy. Overall, the reasoning accuracy increases when the memory rate increases, but is less affected by the search step length. This suggests that solving mathematical problems benefits more from stable accumulation of prior steps and residual refinement (i.e., memory rate). A stronger memory allows the model to preserve intermediate reasoning steps for better reasoning. Figure 6(b) shows the opposite trend. AQUA is less sensitive overall, but is more affected by the search step size. This suggests that common-sense QA relies more on adaptive correction than long-term memory, so flexible updates in latent space have a greater impact than memory accumulation.
@@
Q6: Generalization Across Backbones
To answer Q6, we evaluate the generalization of our framework on different LLM backbones across architectures and sizes: GPT-2 base (117M), Qwen-2.5 1.5B, and LLaMA-3.2 1B models.
Table 2 shows that our method improves accuracy over the baseline on all three backbones. For example, on Qwen and LLaMA, the base accuracy is 42.29% and 39.30%, and our method improves it to 43.04% and 41.10%, with gains of +1.77% and +4.58%, respectively. This confirms that our latent refinement strategy can work across different model architectures and sizes.
The method is simple and general. It does not require any changes to the model architecture and applies equally to small and medium models. It also adds little overhead. Training time and memory increase with model size as expected, but inference remains efficient — all models complete in under 7 minutes and under 24 GB of memory.
Our approach is especially practical for frozen or resource-limited setups. It can be used on top of any existing language model to improve reasoning, without re-training or modifying model weights.
@@
Study of Token Efficiency
Latent reasoning avoids generating intermediate natural language steps (i.e., thoughts), thus, reduce output token number. We compare the average number of generated tokens between Chain-of-Thought (CoT) prompting and our latent reasoning method on the MathQA and AQUA datasets.
@@
Table 3 shows that latent method reduces token usage by over 92% on both datasets. This is because latent steps are performed in a latent space and don’t produce textual outputs at each reasoning step. These results suggest that latent reasoning is not only more compact but also more cost-efficient during inference. This makes it particularly suitable for deployment in resource-constrained settings or large-scale usage.
Related Work
Chain of Thought Reasoning